A local government operates on a calendar-year basis. Prepare journal entries to record the following transactions and events for calendar year 2018.

1. On February 1, 2018, borrowed $400,000 on tax anticipation notes (TANs). The TANs will be repaid with 1.0 percent interest on January 31, 2019.

2. To prepare for issuing financial statements for 2018, accrue interest on the TANs through December 31, 2018.  

3. Invested $100,000 in a certificate of deposit (CD) on April 1, 2018. The CD, which pays interest of 0.8 percent, will mature on September 30, 2018.

4. The CD matured on September 30, 2018.
Answer #1
Date Account Titles Debit Credit
1 Feb-01 Cash $       400,000
      Tax anticipation notes payable $         400,000
(Short term borrowings made)
2 Dec-31 Expenditures - Interest $            3,667 =400000*1%*11/12
       Accrued Interest Payable $             3,667
(Interest accrued for 11 months)
3 Apr-01 Investments $       100,000
       Cash $         100,000
(Investment made in CD)
4 Sep-30 Cash $       100,400 =100000+400
      Investments $         100,000
      Revenues - Interest Income $                 400 =100000*0.8%*6/12
(CD matured)